Ask your own question, for FREE!
Mathematics 7 Online
OpenStudy (anonymous):

A drawer contains four red socks, three white socks, and three blue socks Without looking, you select a sock at random, replace it, and select a secone sock at random. what is the probability that the first sock is blue and the sock sock is red.

OpenStudy (anonymous):

what is the probability the first sock is blue?

OpenStudy (anonymous):

1. 3/5 2. 7/20 3.3/25 4. 7/100

OpenStudy (anonymous):

3/10*4/10=12/100, so answer 3. 3/25

OpenStudy (anonymous):

Thank you so much!

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
15 hours ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
18 hours ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Arizona: does anyone know the percentage rate of 30,000 and 40,000
1 week ago 0 Replies 0 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!